FERMI Machine Status

Status

Run 32
Shift Users Shifts / TeraFERMI, EIS-TIMEX
Beam to Gun
Next users dedicated injection
Last users dedicated injection 2018-02-01
Control Room Fermi (tel. 8800)
Info updated at

Linac / e-beam

RF plants ----
E-beam Energy 1.453 GeV
Bunch Charge 1 pC

Seed laser

Status OPEN
SLU Status laser ON
Pulse Energy 31.8 µJ
Wavelength 260.14 nm

FEL-2 I0 monitor charts (30 minutes)

FEL-2

1st stage 2nd stage
Harmonic 12 5
Wavelength 21.68 nm 4.34 nm
Polarization Right Circular Left Circular

FEL-2 I0 monitor

1st stage 2nd stage
Intensity ---- -0.1 µJ
RMS ---- 375 %

Spectrometer

Wavelength 4.32 nm
Bandwidth 0.087 nm
Last Updated on Friday, 21 August 2015 15:46